OpenStudy (anonymous):

22. A ball is thrown upward with an initial velocity of 4 ft/sec from a height of 10 feet. What is the maximum height that the ball will reach? @calclover

OpenStudy (anonymous):

@undeadknight26

undeadknight26 (undeadknight26):

._. Your making me do other peoples work?!

undeadknight26 (undeadknight26):

Whats the quadratic formula @Shorty1234 ?

undeadknight26 (undeadknight26):

I g2g rn hun sorry </3 This may help you though: https://answers.yahoo.com/question/index?qid=20100801171808AAX9pVD

undeadknight26 (undeadknight26):

@confluxepic this is all you bro

OpenStudy (anonymous):

Do you know this formula? \[s=\frac{ v_{f}^{2}-v_{i}^{2} }{ 2a }\]

OpenStudy (anonymous):

What is \[s _{0}\] I mean what does it mean?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

Sorry, having connection issues.

OpenStudy (anonymous):

This formula is even simpler: H = (initial velocity)^2 / (2*g)

OpenStudy (anonymous):

H is the maximum height.

OpenStudy (anonymous):

I'm wondering, do you get some kind of collection of formulas while doing these kind of problems?

OpenStudy (anonymous):

Anyway the formula was derived from one of the kinematic equations.

OpenStudy (anonymous):

I didn't get one.

OpenStudy (anonymous):

except for the quadratic formula/discriminant

OpenStudy (anonymous):

what's g?

OpenStudy (anonymous):

@aleroth

OpenStudy (anonymous):

g is the acceleration due to gravity, 9.8 m/s^2

OpenStudy (anonymous):

how do i find g??

OpenStudy (anonymous):

It's well known as 9.8

OpenStudy (anonymous):

I dunno, I'm seeing this as a physics problem

OpenStudy (anonymous):

im taking pre-calc so....

OpenStudy (anonymous):

Here's how I'd solve it: H = 4^2 / 2*9.8 = 0.8163 ft (using the formula) 10 + 0.8163 = 10.8163 ft (adding on the 10 we started at)